How would a behaviourist teacher explain 1/2÷1/3? (2)

1
Expert's answer
2022-05-16T17:03:36-0400

Dividing two fractions is the same as multiplying the first fraction by the reciprocal of the second fraction. The first step to dividing fractions is to find the reciprocal (reverse the numerator and denominator) of the second fraction. Next, multiply the two numerators. Then, multiply the two denominators.

1/2 : 1/3 => 1/2 * 3/1 => 1 * 3 / 2 * 1 = 3/2= 1.5

I recommend you watch the video below:

https://youtu.be/tnkPY4UqJ44

Answer: 1.5